Sound Tek Inc. manufactures electronic stereo equipment. The manufacturing process includes printed circuit (PC) board assembly, final assembly, testing, and shipping. In the PC board assembly operation, a number of individuals are responsible for assembling electronic components into printed circuit boards. Each operator is responsible for soldering components according to a given set of instructions. Operators work on batches of 50 printed circuit boards. Each board requires 4 minutes of board assembly time. After each batch is completed, the operator moves the assembled boards to the final assembly area. This move takes 10 minutes to complete.

The final assembly for each stereo unit requires 20 minutes and is also done in batches of 50 units. A batch of 50 stereos is moved into the test building, which is across the street. The move takes 20 minutes. Before conducting the test, the test equipment must be set up for the particular stereo model. The test setup requires 30 minutes. The units wait while the setup is performed. In the final test, the 50-unit batch is tested one at a time. Each test requires 6 minutes. The completed batch, after all testing, is sent to shipping for packaging and final shipment to customers. A complete batch of 50 units is sent from testing to shipping. The Shipping Department is located next to testing. Thus, there is no move time between these two operations. Packaging and labeling requires 7 minutes per unit.

1. Determine the amount of value-added and non-value-added lead time and the value-added ratio in this process for an average stereo unit in a batch of 50 units. Categorize the non-value-added time into wait and move time. Round the percentage to one decimal place.

Value-added lead time min.
Non-value-added lead time:
Wait time lead time min.
Move time lead time min.
Total non-value-added lead time min.
Total lead time min.
Value-added ratio (as a percent) %
Answer #1

Answer:

Minutes Minutes
Value-added lead time 37 a
Non-value-added lead time:
Wait time lead time 1843 b
Move time lead time (10+20) 30 c
Total non-value-added lead time 1873 d=b+c
Total lead time 1910 e=a+d
Value-added ratio (as a percent) 1.9% f=(a/e) x 100 %
a Value-added lead time
Particulars Minutes
PC board assembly 4
Final assembly 20
Testing 6
Packaging and labeling 7
Total Value added time 37
b Wait time lead time
Particulars Minutes
Wait time for Non added value 1,813
[value added time x (batch size -1)]
[ 37 x (50-1)]
Add: Test set-up time 30
Wait time 1,843